OpenStudy (anonymous):

Write the equation of the line that passes through (–2, 6) and (2, 14) in slope-intercept form. y = 2x – 2 y = 2x + 10 y = 0.5x + 7 y = 0.5x – 4

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@BweadedChicken

OpenStudy (anonymous):

y2 - y1 / x2 - x1 <-- first formula, put everything in place then let me see

OpenStudy (anonymous):

14-6 \2-(-2)

OpenStudy (anonymous):

alright, so thats 8/4 which is 2, that was the formula for the slope

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

m = 2 now, cause were using M as a symbol of slope... pick an equation, -2, 6 or 2, 14?

OpenStudy (anonymous):

-2,6

OpenStudy (anonymous):

y - y1 = m(x - x1) plug everything in

OpenStudy (anonymous):

y-6=2(x-(2))

OpenStudy (anonymous):

what would y and x be to plug in

OpenStudy (anonymous):

y and x aint plug ins, theyre just y and x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

id give u the answer, but "Im not allowed to" xD, but distribute the 2 to the ()'s

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

oh wait no

OpenStudy (anonymous):

no, its B

OpenStudy (anonymous):

and u get 10 from the distribute right?

OpenStudy (anonymous):

can i ask one more? :|

OpenStudy (anonymous):

holy hell, my computer lagged, but yeah

OpenStudy (anonymous):

Write the equation of the line that passes through (3, –2) and has a slope of 4 in point-slope form. y + 2 = 4(x – 3) y – 3 = 4(x + 2) x – 3 = 4(y + 2) x + 2 = 4(y – 3)

OpenStudy (anonymous):

alright, we can skip the first part which was y2 - y1 / x2 - x1 cause they already gave u the slope...

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[y - (-2) = 4(x - 3)\]

OpenStudy (anonymous):

distribute the 4 to (x-3)

OpenStudy (anonymous):

oh wait, i already gave u the answer xD

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

oh lol i was about to go through lol thx :)

OpenStudy (anonymous):

lol sorry, and np

OpenStudy (anonymous):

Anne has a total of 3,500 golf balls at home. She loses some golf balls when she goes out on the driving range daily. The number of golf balls, h(x), can be represented by the function h(x) = –150x + 3,500, where x is the number of days she plays golf. Suppose 3,500 changed to 4,000. What does this mean in the context of the problem? Changing 3,500 to 4,000 means the initial number of golf balls decreased. Changing 3,500 to 4,000 means the initial number of golf balls increased. Changing 3,500 to 4,000 means the daily number of golf balls lost decreased. Changing 3,500 to 4,000 means the daily number of golf balls lost increased.

OpenStudy (anonymous):

and that would be D right?

OpenStudy (anonymous):

this is the last i promise haha

OpenStudy (anonymous):

Im not 100% sure cause it says 3500 is how much she has at home, so in a way its increased, but also decreased cause she can lose em all

OpenStudy (anonymous):

im thinking it would be B

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

because its the start off amount, the -150x is how much she loses a day

OpenStudy (anonymous):

yea so thats why I'm lost cuz in a way its decreasing and increasing

OpenStudy (anonymous):

its alright, ask me as much questions as u want any time im online, im usually only chatting anyways, but i gotta head out, cya!

OpenStudy (anonymous):

ight thx again

OpenStudy (anonymous):

thx bro i got a 100%